The proper operation of typical home appliances requires voltage levels that do not vary much. A sample of voltage levels for 10 days was recorded and it was found that the standard deviation of the sample was 0.23 volt. Construct a 90% confidence interval estimate of the standard deviation of all voltage levels. Write a sentence summarizing your results.
Answer #1

Solution :

Given that,

s = 0.23

s2 = 0.0529

n = 10

Degrees of freedom = df = n - 1 = 10 - 1 = 9

At 90% confidence level the \chi 2 value is ,

\alpha = 1 - 90% = 1 - 0.90 = 0.10

\alpha / 2 = 0.10 / 2 = 0.05

1 -  \alpha / 2 = 1 - 0.05 = 0.95

\chi2L = \chi 2\alpha/2,df = 16.919

\chi2R = \chi 21 - \alpha/2,df = 3.325

The 95% confidence interval for \sigma is,

\sqrt(n - 1)s2 / \chi 2\alpha/2 < \sigma < \sqrt (n - 1)s2 / \chi 21 - \alpha/2

\sqrt9 * 0.0529 / 16.919 < \sigma < \sqrt 9 * 0.0529 / 3.325

0.17 < \sigma < 0.38

( 0.17, 0.38)

The 95% confidence interval for \sigma is, ( 0.17, 0.38) .